# 3D Cell Oversegmentation Correction via Geo-Wasserstein Divergence

### v3

@@ Abstract
3D cell segmentation methods are often hindered by oversegmentation, where a single cell is incorrectly split into multiple fragments. This degrades the final segmentation quality and is notoriously difficult to resolve, as oversegmentation errors often resemble natural gaps between adjacent cells. Our work makes two key contributions. First, for 3D cell segmentation, we are the first work to formulate oversegmentation as a concrete problem and propose a geometric framework to identify and correct these errors. Our approach builds a pre-trained classifier using both 2D geometric and 3D topological features extracted from flawed 3D segmentation results. Second, we introduce a novel metric, Geo-Wasserstein divergence, to quantify changes in 2D geometries. This captures the evolving trends of cell mask shape in a geometry-aware manner. We validate our method through extensive experiments on in-domain plant datasets, including both synthesized and real oversegmented cases, as well as on out-of-domain animal datasets to demonstrate transfer learning performance. An ablation study further highlights the contribution of the Geo-Wasserstein divergence. A clear pipeline is provided for end-users to build pre-trained models to any labeled dataset.
@@ 1 Introduction
3D cell segmentation aims to reconstruct complete 3D cellular structures from sequential 2D microscopy images, a crucial step for downstream analysis of cellular characteristics such as morphology, cell type classification, and volumetric properties [4, 34]. However, noises in raw microscopy images can propagate into the final 3D segmentation, resulting in oversegmentation. Shown in Figure 1, this challenge occurs when individual cells are mistakenly split into multiple segments, which largely downgrades the quality of 3D segmentation results produced the current state-of-the-art methods.
@@
(ii) We introduce a novel geometry-aware metric, Geo-Wasserstein divergence, that leverages earthmover’s distance to quantify the divergence between 2D geometric shapes of cell masks. We further highlight its potential application to broader scenarios in general segmentation task for the similar trait of temporal consistency.
@@ 2 Preliminaries and Related Works
The development of cell segmentation has been greatly facilitated by deep learning architectures such as U-Net [22], which have given rise to several state-of-the-art 2D segmentation methods, including CellPose2D [19, 27, 28], Mesmer [16], and StarDist [25]. Building on these advances, 3D segmentation methods can be broadly categorized into two types: (i) 2D-based methods, which typically rely on unsupervised or weakly supervised 2D segmentation followed by subsequent 3D reconstruction, thereby benefiting from reduced complexity and strong generalizability; and (ii) Non-2D-based methods, which directly reconstruct 3D structures from raw image stacks, generally requiring fully supervised training on large labeled datasets. Within it, 2.5D-based approaches have emerged, which integrates the 2D pre-segmented information into the final 3D contextual analysis. Such hybrid framework avoids the need for large-scale 3D pretraining while leveraging the strong generalizability of 2D-based methods. CellPose [28] (2.5D-based), CellStitch [17] (2D-based), and PlantSeg [35] (non-2D-based) represent the state-of-the-art approaches in their respective categories. However, 3D-based or deep learning-based methods [8, 31, 32] such as PlantSeg often suffers from limited generalizability due to its reliance on supervised pretraining with large labeled datasets. We review the method above and provide a technical review for more 2D segmentation backbones [22, 15, 21, 7, 13, 36] and 3D segmentation approaches [30, 1, 33, 29] in Appendix A.
@@ 2.2 Optimal Transport
Optimal transport (OT) has been applied to mutiple biological areas, including single-cell studies [24], drug perturbations [5, 10, 11, 9], and cell alignments [12]. In our framework, OT is utilized to quantify the geometric change between 2D cell masks from adjacent layers. We introduce OT formulation along with earthmover’s distance (EMD) below.
@@
Formulation (2.2), known as the Kantorovich formulation of optimal transport, generalizes the Monge formulation [18] by relaxing the requirement of a direct one-to-one mapping between source and target elements. For two discrete distributions $X=\{x_{i}\}_{i=1}^{n}$ and $Y=\{y_{j}\}_{j=1}^{m}$, the earthmover’s distance is defined as the minimum cost required to transform $X$ into $Y$:
@@
The EMD provides a measurement of the distance between two distributions based on the minimal work required to transform one distribution into the other. Utilizing the EMD, we introduce a novel approach to quantify the divergence between two 2D geometries: Geo-Wasserstein divergence. By approximating 2D geometries into discrete distributions of mass (see §3.2), EMD allows us to calculate the minimal transportation cost required to transform one geometry into the other. When cell masks get larger, computing the EMD becomes computationally expensive. Still, sliced Wasserstein distance [3] can be used a computationally efficient approach to estimate EMD. Typically, increasing the number of projections in sliced Wasserstein method would bring better approximation to the true EMD.
@@ 3.2 Method Pipeline
Through extensive experiments on both plant and animal cells, we validate this assumption in the context of 3D cellular image segmentation, which forms the basis for the subsequent geometric feature extraction framework using Geo-Wasserstein divergence.
@@
We propose a straightforward method from 3D topological perspective to distinguish oversegmented cell pairs from those separated by natural gaps. As illustrated in Figure 4, a key pattern observed across multiple datasets is that oversegmented cases typically exhibit a smooth shape transition at the gap layer, often following either a linear or circular progression. These patterns have biological relevance: circular shapes—wider in the middle and narrower at the ends—are characteristic of mitotic cells undergoing “circular rounding” [23, 6], while gradually tapering linear shapes are typical of structural cells such as palisade cells [20]. We quantify such linearity and circularity by analyzing the overlapping area between masks at adjacent layers. For cells exhibiting a linear pattern, the overlap area should change in a monotonically linear trend. While for circular patterns, the overlap area is expected to follow a strictly quadratic trend, characterized by a monotonous increase followed by a decrease. As shown in the Table 1, we evaluate this strict pattern across cells in the labeled datasets:
@@
For final oversegmented cell pairs, we recover the mis-segmented 2D mask between them (at layer $i$) by interpolating between the cell mask boundaries at layers i1i-1 and $i+1$, leveraging the interpolation methods [17, 26]. The corrected 2D segmentation masks are then used to reconstruct the updated 3D segmentation. The technical details of cross-layer interpolation are provided in Appendix B.1.
@@ 4 Experiments
We evaluate our method on two types of datasets: publicly available, labeled plant cell datasets, and private, unlabeled animal cell datasets. Specifically, the animal datasets lack ground-truth annotation. In total, we use six plant cell datasets [2], each consisting of 100 image stacks, and an animal cell datasets with 11 image stacks. Notably, each animal image stack is a large volumetric dataset containing approximately 300 image layers, with the maximum $\mathrm{X}\times\mathrm{Y}$ plane size reaching up to $2008\times 2008$ pixels. After the initial 3D segmentation, each image stack contains approximately 2,000 to 4,000 cells. Note that due to the scarcity of high-quality and labeled 3D volumetric image stacks, plant datasets from [2] are commonly used as a standard benchmark for 3D cell segmentation training, as adopted by state-of-the-art methods [17, 14, 35, 31]. To demonstrate the effectiveness under transfer learning, we additionally evaluated on our private animal cell datasets without re-training, showing its applicability across different cell types in organ-level image stacks. Further dataset details are provided in Appendix D.
@@ 5 Conclusion
We present a geometric framework for 3D cell oversegmentation, an important yet unresolved challenge that degrades the quality of 3D segmentation results. Beyond its impact on 3D cell segmentation, the proposed Geo-Wasserstein divergence is broadly applicable to segmentation problems, especially video segmentation, where temporal consistency mirrors the slice-to-slice geometric consistency exploited in 3D cell stacks: in both settings, the key question is whether regions in neighboring frames (time) or slices (depth) correspond to the same object. We hope this framework facilitates adoption by biological end-users and inspires the computer vision and geometry processing community to build upon it.
Due to space constraints, we provide additional visualizations across datasets in Appendix C of the supplementary material. Further methodological and dataset details appear in Appendix B and Appendix D, respectively.
Acknowledgement
We acknowledge computing resources from Columbia University’s Shared Research Computing Facility project, which is supported by NIH Research Facility Improvement Grant 1G20RR030893-01, and associated funds from the New York State Empire State Development, Division of Science Technology and Innovation (NYSTAR) Contract C090171, both awarded April 15, 2010.
PC and YL would also like to express our appreciation to Andrew Blumberg, Abigail Hickok, and Yinuo Jin for their valuable feedback and discussions.
@@ Appendix A Further Related Works
U-Net [22] is a fully convolutional neural network specifically developed for biomedical image segmentation. It employs an encoder-decoder structure with skip connections that preserve spatial information, making it ideal for cell and tissue segmentation. Hover-Net [15] combines semantic segmentation with instance-level clustering via predicted horizontal and vertical distance maps (hover maps). Hover-Net is particularly effective for segmenting densely packed cell nuclei in pathology slides. Micro-Net [21] is a lightweight CNN architecture specifically tailored for segmenting cells in resource-limited environments. It leverages depth-wise separable convolutions and inverted residuals to achieve efficient segmentation with fewer parameters. DCAN [7] simultaneously predicts pixel-wise object masks and object contours using a dual-branch CNN architecture. Effective for precisely delineating individual cells by explicitly modeling boundaries. Omni-Seg [13] uses an ensemble of CNN models trained on multiple image types to generalize across a wide range of biological imaging modalities without retraining. It applies domain adaptation and multi-task learning techniques. NuSeT [36] is a CNN model tailored specifically for nuclear segmentation in fluorescence microscopy images, combining U-Net structure with a tailored loss function to enhance segmentation accuracy in noisy conditions.
@@
3DCellSeg [30] is a two-stage deep-learning pipeline designed for dense 3D cell segmentation in membrane-stained images. The pipeline is robust with only one main hyperparameter, avoiding extensive manual tuning and generalizing across different datasets and cell shapes. But still, as a supervised 3D-based method, 3DCellSeg still requires annotated training data for new cell types or imaging conditions. Its performance can degrade if applied to modalities it wasn’t trained on, necessitating retraining. Additionally, like many 3D-based methods, processing very large volumetric images can be computationally intensive comparing to 2D-based methods.
CellSeg3D [1] introduces a self-supervised 3D cell segmentation approach that eliminates the need for manual annotations. It employs a WNet3D architecture—comprising two 3D U-Nets connected sequentially—to perform segmentation by optimizing a soft normalized cuts loss directly on raw image volumes. However, as a self-supervised method, it relies on inherent image features, such as brightness differences, to delineate cells. This reliance may limit its ability to generalize effectively to end-user datasets, which often exhibit lower quality. While the performance is well on colorful fluorescent images, it may not be suitable for application in our generalized setting.
StarDist3D [33] is an instance segmentation method that represents cells (typically for nuclei) as star-convex shapes. A CNN predicts, for each pixel/voxel, the distances to the object boundary along a fixed set of radial rays, as well as an object probability score. StarDist assumes objects are approximately star-convex, so it may struggle with cells that have extremely irregular or concave shapes not well represented by a single star-shaped model. Another practical limitation is the need for training data for each new application. Due to the lack of large 3D training datasets, currently has no specific pretrained 3D StarDist models are publicly available. Still, StarDist3D achieves good performance on its specific cell types.
Go-Nuclear [29] is a recently introduced toolkit for 3D nucleus segmentation in tissue and organ datasets. Rather than a single algorithm, it encompasses a pipeline for generating training data and iteratively training models using multiple algorithms. The large curated dataset is used to train and fine-tune popular segmentation frameworks (CellPose, PlantSeg, StarDist) for nuclei in diverse organs. Yet, Go-Nuclear models are specialized for nuclear segmentation and rely on having a clearly labeled nucleus channel.

### v4

@@ Abstract
3D cell segmentation methods are often hindered by oversegmentation, where a single cell is incorrectly split into multiple fragments. This degrades the final segmentation quality and is notoriously difficult to resolve, as oversegmentation errors often resemble natural gaps between adjacent cells. Our work makes two key contributions. First, we formulate 3D cell oversegmentation as a concrete learning problem and propose a geometry-aware framework to identify and correct these errors. Our approach builds a pre-trained classifier using both 2D geometric and 3D topological features extracted from flawed 3D segmentation results. Second, we introduce a novel metric, Geo-Wasserstein divergence, to quantify changes in 2D geometries. This captures the evolving trends of cell mask shape in a geometry-aware manner. We validate our method through extensive experiments on in-domain plant datasets, including both synthesized and real oversegmented cases, as well as on out-of-domain animal datasets to demonstrate transfer learning performance. An ablation study further highlights the contribution of the Geo-Wasserstein divergence. A clear pipeline is provided for end-users to build pre-trained models to any labeled dataset.
@@ 1 Introduction
3D cell segmentation aims to reconstruct complete 3D cellular structures from sequential 2D microscopy images, a crucial step for downstream analysis of cellular characteristics such as morphology, cell type classification, and volumetric properties [boutros2015, williams2022]. However, noises in raw microscopy images can propagate into the final 3D segmentation, resulting in oversegmentation. Shown in Figure 1, this challenge occurs when individual cells are mistakenly split into multiple segments, which largely downgrades the quality of 3D segmentation results produced the current state-of-the-art methods.
@@
(ii) We introduce a novel geometry-aware shape descriptor, Geo-Wasserstein divergence, that leverages earthmover’s distance to quantify the divergence between 2D geometric shapes of cell masks. We further highlight its potential application to broader scenarios in general segmentation task for the similar trait of temporal consistency.
@@ 2 Preliminaries and Related Works
The development of cell segmentation has been greatly facilitated by deep learning architectures such as U-Net [ronneberger2015unet], which have given rise to several state-of-the-art 2D segmentation methods, including CellPose2D [cellpose2.0, cellpose3.0, cellpose1.0], Mesmer [greenwald2022wholecell_segmentation], and StarDist [schmidt2018starconvex]. Building on these advances, 3D segmentation methods can be broadly categorized into two types: (i) 2D-based methods, which typically rely on unsupervised or weakly supervised 2D segmentation followed by subsequent 3D reconstruction, thereby benefiting from reduced complexity and strong generalizability; and (ii) Non-2D-based methods, which directly reconstruct 3D structures from raw image stacks, generally requiring fully supervised training on large labeled datasets. Within it, 2.5D-based approaches have emerged, which integrates the 2D pre-segmented information into the final 3D contextual analysis. Such hybrid framework avoids the need for large-scale 3D pretraining while leveraging the strong generalizability of 2D-based methods. CellPose [cellpose1.0] (2.5D-based), CellStitch [Liu2023Cellstitch] (2D-based), and PlantSeg [plantseg] (non-2D-based) represent the state-of-the-art approaches in their respective categories. However, 3D-based or deep learning-based methods [DL1, DL2, DL3] such as PlantSeg often suffers from limited generalizability due to its reliance on supervised pretraining with large labeled datasets. Technical reviews and baseline choice explanations for more 3D segmentation approaches [wang2022novel, achard2024cellseg3d, DL3, vijayan2024deep], 2D segmentation backbones [ronneberger2015unet, Graham2019HoverNet, raza2019micro, Chen_2016_CVPR, deng2023omni, yang2020nuset], and geometry-aware segmentation methods [hu2019topology, pmlr-v202-stucki23a, lux2025topographefficientgraphbasedframework, hu2021topology, Shit_2021] are provided in Appendix A.
@@ 2.2 Optimal Transport
Optimal transport (OT) has been applied to mutiple biological areas, including single-cell studies [schiebinger2019optimal], drug perturbations [bunne2023cellot, chen2024sicnn, chen2025displacement], and cell alignments [demetci2022scot]. In our framework, OT is utilized to quantify the geometric change between 2D cell masks from adjacent layers. We introduce OT formulation along with earthmover’s distance (EMD) below.
@@
Formulation (2.2), known as the Kantorovich formulation of optimal transport, generalizes the Monge formulation [Monge1781] by relaxing the requirement of a direct one-to-one mapping between source and target elements. For two discrete distributions $X=\{x_{i}\}_{i=1}^{n}$ and $Y=\{y_{j}\}_{j=1}^{m}$, the earthmover’s distance is defined as the minimum cost required to transform $X$ into $Y$:
@@
The EMD provides a measurement of the distance between two distributions based on the minimal work required to transform one distribution into the other. Utilizing the EMD, we introduce a novel approach to quantify the divergence between two 2D geometries: Geo-Wasserstein divergence. By approximating 2D geometries into discrete distributions of mass (see § 3.2), EMD allows us to calculate the minimal transportation cost required to transform one geometry into the other. When cell masks get larger, computing the EMD becomes computationally expensive. Still, sliced Wasserstein distance [bonneel2015sliced] can be used a computationally efficient approach to estimate EMD. Typically, increasing the number of projections in sliced Wasserstein method would bring better approximation to the true EMD.
@@ 3.2 Method Pipeline
Through extensive experiments on plant cells, we validate this assumption in 3D cellular image segmentation. We further show that this pattern generalizes beyond the training distribution: on eleven animal cell stacks, which lack cell walls and therefore exhibit greater irregularity than plant cells, our approach remains effective.
@@
We propose a straightforward method from 3D topological perspective to distinguish oversegmented cell pairs from those separated by natural gaps. As illustrated in Figure 4, a key pattern observed across multiple datasets is that oversegmented cases typically exhibit a smooth shape transition at the gap layer, often following either a linear or circular progression. These patterns have biological relevance: circular shapes—wider in the middle and narrower at the ends—are characteristic of mitotic cells undergoing “circular rounding” [cr2, cr1], while gradually tapering linear shapes are typical of structural cells such as palisade cells [pallardy2008physiology]. We quantify such linearity and circularity by analyzing the overlapping area between masks at adjacent layers. For cells exhibiting a linear pattern, the overlap area should change in a monotonically linear trend. While for circular patterns, the overlap area is expected to follow a strictly quadratic trend, characterized by a monotonous increase followed by a decrease. As shown in the Table 1, we evaluate this strict pattern across cells in the labeled datasets:
@@
For final oversegmented cell pairs, we recover the mis-segmented 2D mask between them (at layer $i$) by interpolating between the cell mask boundaries at layers i1i-1 and $i+1$, leveraging the interpolation methods [Liu2023Cellstitch, Solomon2015]. The corrected 2D segmentation masks are then used to reconstruct the updated 3D segmentation. The technical details of cross-layer interpolation are provided in Appendix B.1.
@@ 4 Experiments
We evaluate our method on two types of datasets: publicly available, labeled plant cell datasets, and private, unlabeled animal cell datasets. Specifically, the animal datasets lack ground-truth annotation. In total, we use six plant cell datasets [bassel2019_dataset], each consisting of 100 image stacks, and an animal cell datasets with 11 image stacks. Notably, each animal image stack is a large volumetric dataset containing approximately 300 image layers, with the maximum $\mathrm{X}\times\mathrm{Y}$ plane size reaching up to $2008\times 2008$ pixels. After the initial 3D segmentation, each image stack contains approximately 2,000 to 4,000 cells. Note that due to the scarcity of high-quality and labeled 3D volumetric image stacks, plant datasets from [bassel2019_dataset] are commonly used as a standard benchmark for 3D cell segmentation training, as adopted by state-of-the-art methods [Liu2023Cellstitch, eschweiler2022robust3dcellsegmentation, plantseg, DL2]. To demonstrate the effectiveness under transfer learning, we additionally evaluated on our private animal cell datasets without re-training, showing its applicability across different cell types in organ-level image stacks. Further dataset details are provided in Appendix D.
@@ 5 Conclusion
We present a geometric framework for 3D cell oversegmentation, an important yet unresolved challenge that degrades the quality of 3D segmentation results. Beyond its impact on 3D cell segmentation, the proposed Geo-Wasserstein divergence is broadly applicable to segmentation problems, especially video segmentation, where temporal consistency mirrors the slice-to-slice geometric consistency exploited in 3D cell stacks: in both settings, the key question is whether regions in neighboring frames (time) or slices (depth) correspond to the same object. We hope this framework facilitates adoption by end-users and inspires the vision and geometry processing community to build upon it.
Acknowledgment. We acknowledge computing resources from Columbia University’s Shared Research Computing Facility project, which is supported by NIH Research Facility Improvement Grant 1G20RR030893-01, and associated funds from the New York State Empire State Development, Division of Science Technology and Innovation (NYSTAR) Contract C090171, both awarded April 15, 2010.
@@ Appendix A Further Related Works
U-Net [ronneberger2015unet] is a fully convolutional neural network specifically developed for biomedical image segmentation. It employs an encoder-decoder structure with skip connections that preserve spatial information, making it ideal for cell and tissue segmentation. Hover-Net [Graham2019HoverNet] combines semantic segmentation with instance-level clustering via predicted horizontal and vertical distance maps (hover maps). Hover-Net is particularly effective for segmenting densely packed cell nuclei in pathology slides. Micro-Net [raza2019micro] is a lightweight CNN architecture specifically tailored for segmenting cells in resource-limited environments. It leverages depth-wise separable convolutions and inverted residuals to achieve efficient segmentation with fewer parameters. DCAN [Chen_2016_CVPR] simultaneously predicts pixel-wise object masks and object contours using a dual-branch CNN architecture. Effective for precisely delineating individual cells by explicitly modeling boundaries. Omni-Seg [deng2023omni] uses an ensemble of CNN models trained on multiple image types to generalize across a wide range of biological imaging modalities without retraining. It applies domain adaptation and multi-task learning techniques. NuSeT [yang2020nuset] is a CNN model tailored specifically for nuclear segmentation in fluorescence microscopy images, combining U-Net structure with a tailored loss function to enhance segmentation accuracy in noisy conditions.
@@
3DCellSeg [wang2022novel] is a two-stage deep-learning pipeline designed for dense 3D cell segmentation in membrane-stained images. The pipeline is robust with only one main hyperparameter, avoiding extensive manual tuning and generalizing across different datasets and cell shapes. But still, as a supervised 3D-based method, 3DCellSeg still requires annotated training data for new cell types or imaging conditions. Its performance can degrade if applied to modalities it wasn’t trained on, necessitating retraining. Additionally, like many 3D-based methods, processing very large volumetric images can be computationally intensive comparing to 2D-based methods.
CellSeg3D [achard2024cellseg3d] introduces a self-supervised 3D cell segmentation approach that eliminates the need for manual annotations. It employs a WNet3D architecture—comprising two 3D U-Nets connected sequentially—to perform segmentation by optimizing a soft normalized cuts loss directly on raw image volumes. However, as a self-supervised method, it relies on inherent image features, such as brightness differences, to delineate cells. This reliance may limit its ability to generalize effectively to end-user datasets, which often exhibit lower quality. While the performance is well on colorful fluorescent images, it may not be suitable for application in our generalized setting.
StarDist3D [weigert2020stardist3d] is an instance segmentation method that represents cells (typically for nuclei) as star-convex shapes. A CNN predicts, for each pixel/voxel, the distances to the object boundary along a fixed set of radial rays, as well as an object probability score. StarDist assumes objects are approximately star-convex, so it may struggle with cells that have extremely irregular or concave shapes not well represented by a single star-shaped model. Another practical limitation is the need for training data for each new application. Due to the lack of large 3D training datasets, currently has no specific pretrained 3D StarDist models are publicly available. Still, StarDist3D achieves good performance on its specific cell types.
Go-Nuclear [vijayan2024deep] is a recently introduced toolkit for 3D nucleus segmentation in tissue and organ datasets. Rather than a single algorithm, it encompasses a pipeline for generating training data and iteratively training models using multiple algorithms. The large curated dataset is used to train and fine-tune popular segmentation frameworks (CellPose, PlantSeg, StarDist) for nuclei in diverse organs. Yet, Go-Nuclear models are specialized for nuclear segmentation and rely on having a clearly labeled nucleus channel.
@@
Geometry-aware Segmentation Methods.
We further introduce several geometry- and topology-aware segmentation methods. Hu et al. [hu2019topology, hu2021topology] propose loss functions and training schemes that encourage a segmentation network to preserve topological structures during end-to-end training from raw images. lux2025topographefficientgraphbasedframework define a topology-preserving loss framework based on a component graph for the joint ground-truth/prediction topology, again used during network training. pmlr-v202-stucki23a introduce Betti matching, a spatially accurate loss for segmentation, by computing induced matchings between the persistence barcodes of prediction and ground truth. Shit_2021 propose centerline Dice, a similarity measure based on the intersection of segmentation masks with their skeletons, designed for tubular network-like structures, which is beyond the anisotropic 3D cellular segmentation studied in our work.
In other words, existing geometry-aware methods primarily address how to train a segmentation model, rather than how to perform a specific post-processing step on the output of such a model. These approaches are therefore complementary, not competing, to ours. All of them focus on loss functions for training a segmentation network from scratch or for fine-tuning it; their output is the segmentation itself. Moreover, these works mainly preserve global topology of one or a few foreground classes, whereas cell oversegmentation is an instance-level phenomenon. Frequently, Betti numbers and global connectivity are unchanged whether a single cell is split into two labels or kept as one, so these losses are blind to the specific oversegmentation errors that occur in a post-hoc manner. Adapting the ideas from the above methods to our setting would require substantial re-design to make them suitable as independent post-processing procedures for cellular segmentation. Therefore, we choose the most widely-used, state-of-the-art baselines and validate whether our method can further refine the segmentation quality over these methods. We hope that our framework will inspire follow-up work to adapt geometry-aware losses from prior segmentation methods into post-hoc processing methods.